# Your Classifier Can Do More: Towards Bridging the Gaps in Classification, Robustness, and Generation

### v1

@@ 3.3 Energy-based Joint Adversarial Distribution Training
where $\mathcal{D}$ denotes training data. The inner maximization problem aims to move the adversarial examples out of the high-density area, which corresponds to minimize the joint probability $\log{p_{\theta}}\left(\tilde{\mathbf{x}},y\right)$. To ensure that adversarial examples look more natural, we employ perceptual loss to maintain visual similarity to the clean samples. Specifically, we update SGLD process along the opposite direction of $\Delta_{\mathbf{x}}\log{p_{\theta}}\left(\tilde{\mathbf{x}},y\right)$ to sample the adversarial examples:
$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{c^{2}}{2}\frac{\partial}{\partial\tilde{\mathbf{x}}_{t}}\left[\log{p_{\theta}}\left(\tilde{\mathbf{x}}_{t},y\right)+\lambda d\left(\mathbf{x}_{t},\tilde{\mathbf{x}}_{t}\right)\right],$ (12)
where $\lambda$ is the weight, $d\left(\mathbf{x}_{t},\tilde{\mathbf{x}}_{t}\right)$ denotes the perceptual distance between adversarial and clean samples, which is defined via LPIPSzhang2018unreasonable , which employs a pre-trained network $h(\cdot)$ to compute the feature distances. We begin by extracting activations from $L$ layers of the function $h(\cdot)$ and normalize each across the channel dimension, yielding $\tilde{h}_{l}(x)$ for each layer $l$. Subsequently, each $\tilde{h}_{l}(x)$ is further normalized according to the layer’s size. Using these normalized activations, the distance $d\left(\mathbf{x},\tilde{\mathbf{x}}\right)$ is then defined as follows:
$d(\mathbf{x},\tilde{\mathbf{x}})=\sum_{l}^{L}w_{l}\left\|\frac{\tilde{h}_{l}(\mathbf{x})}{\sqrt{W_{l}H_{l}}}-\frac{\tilde{h}_{l}(\tilde{\mathbf{x}})}{\sqrt{W_{l}H_{l}}}\right\|_{2}^{2},$ (13)
where $W_{l}$ and $H_{l}$ denote the width and height of the activations in layer $l$, $w_{l}$ denotes the weight.
On the other hand, the goal of the outer minimization problem is to find model parameters so that the energy difference between adversarial and clean samples is minimized, i.e. pulling back the adversarial examples back to the high-density area. Unlike existing robust classifiers only maximize the conditional probability $log\ p_{\theta}(y\mid\tilde{\mathbf{x}})$, we capture the full adversarial data distribution during training. Similar for Eq. 9, $\frac{\partial{log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}$ can be approximitated as:
@@
$\frac{\partial{log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\tilde{\mathbf{x}}_{i}|\mathbf{x}_{i}^{+}\right)\right].$ (14)
@@
1 Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M$: sampling iterations; $M_{adv}$: adversarial sampling iterations; $M_{\theta}$: sampling iterations for $\theta$ ;
2 Output: $\{\theta_{1},\dots,\theta_{N}\}$: network weights;
3 Init: randomly initialize $\{\theta_{1},\dots,\theta_{N}\}$;
@@ 8             Obtain ${\mathbf{x}}_{0}$ via random noise;
12      Compute $h_{1}=\frac{\partial logp_{\theta}({\mathbf{x}})}{\partial\theta}$ via Eq. 9;
@@
15             Obtain $\tilde{\mathbf{x}}_{0}$ from ${\mathbf{x}}$ with a perturbation;
@@ 4.1 Experimental Setup
We conduct analysis and comparison experiments on several standard classifier datasets, including CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning . Given that most AT methods utilize WideResNetzagoruyko2016wide architectures, we select the commonly used WRN28-10 as the baseline network. During the entire training process, the learning rate is set to 0.01 and the number of sampling steps is 10. The initialization perturbation for adversarial samples is set to 8/255, and the number of adversarial sampling steps is 5. More details are shown in Appendix B. Note that EB-JDAT is a general optimization framework for JEM that is compatible with various JEM-based modelsgrathwohl2019your ; yang2021jem++ ; yang2023towards ; yang2023m . Therefore, we integrate EB-JDAT with SADAJEMyang2023towards and JEM++yang2021jem++ , two faster and more stable JEMs, to achieve improved performance.
@@ 4.2 Comparison Experiments and Analysis
We compare our approach with a range of AT methods, including five popular ones: PGD-AT mkadry2017towards , TRADES zhang2019theoretically , MART wang2019improving , AWP wu2020adversarial , and LBGAT cui2021learnable , as well as four SOTA methods: LAS-AWP jia2022adversarial , TDAT tong2024taxonomy , AGR-TRADES tong2024balancing , and VART zhao2024delving . All AT methods use the WRN34-10 network. For evaluation, we select the PGD mkadry2017towards and AutoAttack (AA) croce2020reliable . The results on CIFAR-10 and CIFAR-100 are shown in Tab. 2, where EB-JDAT not only maintains the classification accuracy of clean data(92.28% on CIFAR-10; 70.19% on CIFAR-100) but also achieves robustness that even surpasses existing SOTA AT methods(+5.96% on CIFAR-10; +2.02% on CIFAR-100). Thus, our method effectively balances the trade-off between accuracy on clean data and robustness against adversarial attacks. Furthermore, EB-JDAT is compatible with both SADAJEMyang2023towards and JEM++yang2021jem++ , thereby demonstrating the universality of our approach.
Table 2: Camparative robustness (%) analysis with different AT methods.
@@
Standard Training 96.10 0 0 80.73 0 0
PGD-AT 85.17 55.08 51.69 60.93 26.05 24.33
TRADES 85.72 56.10 53.40 57.10 26.96 24.55
MART 85.17 58.56 51.10 57.29 26.03 23.92
AWP 85.57 58.13 53.90 60.38 33.86 28.86
LBGAT 88.22 54.66 52.23 60.64 34.75 29.33
LAS-AWP 87.74 60.16 55.52 64.89 36.36 30.77
TDAT 82.25 56.03 54.06 57.32 33.17 26.61
AGR-TRADES 86.50 52.56 52.56 58.17 28.34 25.84
VART 84.21 58.32 54.24 / / /
EB-JDAT-JEM++ 91.43 65.69 64.34 70.00 37.14 35.21
EB-JDAT-SADAJEM 92.28 66.12 64.76 70.19 38.38 35.63
Comparisons with AT Using Generated Data.
Another strategy for improving robustness using generative models is to incorporate additional training data generated by generative models in AT pang2022robustness ; gowal2021improving . Therefore we compare EB-JDAT with SCORE pang2022robustness , Better DM wang2023better and the method proposed in gowal2021improving . SOCRE and Better DM utilize the SOTA Diffusion models ho2020denoising to generated addition training data, while gowal2021improving generate addition data from a ensemble of DDPM ho2020denoising , GANs goodfellow2014generative and VDVAE childvery generators. All methods are implemented using the WRN28-10 zagoruyko2016wide . The results are presented in Tab. 3. Although EB-JDAT does not rely on additional generated data or increased training epochs, it still achieves competitive performance in terms of clean accuracy and robustness. EB-JDAT even outperforms SCORE-1M, Better DM-1M and gowal2021improving . Unlike discriminative learning approaches pang2022robustness ; gowal2021improving ; wang2023better , which struggles to capture the underlying structure of both clean and adversarial data distributions, EB-JDAT, as a hybrid generative model, provides more accurate gradient estimations of clean/adversarial data densities, leading to substantial improvements in both accuracy and robustness. Notably, this is achieved without requiring additional training data and with minimal computational overhead.
Table 3: Camparative robustness (%) analysis with AT methods that utilize generative models for data augmentation, P denotes EB-JDAT with perceptual loss.
      Data       Generate data       Epoch       Clean       AA
@@
      SCORE pang2022robustness       1M       400       88.10       61.51
      The method proposed by gowal2021improving       100M       2000       87.50       63.38
      Better DM wang2023better       1M       400       91.12       63.35
      Better DM wang2023better       50M       1600       92.27       67.17
      EB-JDAT-SADAJEM(P)       n/a       100       92.28       64.76
      EB-JDAT-SADAJEM       n/a       100       89.79       66.30
@@
      SCORE pang2022robustness       1M       400       62.08       31.40
      Better DM wang2023better       1M       400       68.06       35.65
      Better DM wang2023better       50M       1600       72.58       38.83
      EB-JDAT-SADAJEM(P)       n/a       100       70.19       35.63
      EB-JDAT-SADAJEM       n/a       100       68.35       38.46
Compared with JEMs and Energy-based AT.
In this section, we compare EB-JDAT with other JEMs grathwohl2019your ; yang2021jem++ ; yang2023towards and energy-based AT zhu2021towards ; mirza2024shedding in terms of classification accuracy, robustness and generative performance. Sec. 4.2 presents an evaluation of our method against JEM variants with respect to classification accuracy and generation quality. To assess generation quality, we employ the Fréchet Inception Distance (FID) and Inception Score (IS) as metrics. Compared to JEMs, EB-JDAT maintains comparable accuracy and generative performance, while EB-JDAT-SADAJEM even outperforms JEM(FID-10.98) and JEM++(FID-9.7) in terms of generation. Moreover, EB-JDAT exhibits significantly improved robustness over JEMs, as shown in Fig. 2. Thus, EB-JDAT bridges the gap between classification, robustness, and generation.
Table 4: Performance comparison of different hybrid models on CIFAR-10, IF denotes informative initialization.
@@
JEM grathwohl2019your 92.90 38.40 8.76
JEM++ yang2021jem++ 93.73 37.12 8.29
SADAJEM yang2023towards 96.03 11.40 8.63
@@
JEAT zhu2021towards 85.16 38.24 8.80
WEAT mirza2024shedding (PCA) 83.36 30.74 8.97
WEAT mirza2024shedding (IF) 83.36 177.92 3.50
EB-JDAT-JEM++ 91.43 38.89 7.66
EB-JDAT-SADAJEM 92.28 27.42 8.05
Figure 2: Comparative robustness (%) analysis with different energy-based AT.
Next, we compare EB-JDAT with energy-based AT zhu2021towards ; mirza2024shedding . First, we compare the accuracy and robustness with them, which is reported in Fig. 2. The results demonstrate that our method significantly enhances robustness with minimal accuracy degradation, outperforming existing energy-based AT. Moreover, our method yields superior generative performance compared to energy-based methods, which are shown in Sec. 4.2. EB-JDAT surpasses JEAT and WEAT with an FID reduction of 10.82 and 3.32, respectively. Unlike JEAT zhu2021towards directly introduces adversarial samples into the JEMs training process to model the joint distribution ${p}_{\theta}(\tilde{\mathbf{x}},y)$, we model the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ to capture the distribution of clean and adversarial samples more accurately, thus performing better in accuracy, robustness and generative ability. On top of that, WEATmirza2024shedding reinterprets TRADESzhang2019theoretically as an EBM to model the conditional distribution $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$, which is still a discriminative model. Although mirza2024shedding notes that PCA-based initialization can give WEAT generative capabilities(FID=30.74), but under the same comparison of informative initialization, WEAT shows poor generative performance(FID=177.92). Consequently, EB-JDAT outperforms previous energy-based AT in classification accuracy, robustness, and generative capability.
@@
To further visually explore generative ability of EB-JDAT, we generate images via SGLD sampling with informative initialization for conventional AT-trained models, energy-based AT-models, JEMs, and EB-JDAT in Fig. 3. As demonstrated in Fig. 3, the conventional AT-trained models(a, b) and energy-based AT-models(e) possess limited generative capacity, resulting in difficulties in generating high-quality images. Conversely, EB-JDAT produce higher-resolution images with richer background details, which is more closely matching the features of the original dataset. In addition, EB-JDAT(g, h) surpasses JEM(f) and perform competitively with JEM++ and SADAJEM in generative capacity, producing more diverse and detailed images with only a slight reduction in sharpness, particularly in the ’car’ category.

### v2

@@ 3.3 Energy-based Joint Distribution Adversarial Training
where $\mathcal{D}$ denotes training data. The inner maximization problem aims to move the adversarial examples out of the high-density area, which corresponds to minimize the joint probability $\log{p_{\theta}}\left((\tilde{\mathbf{x}}\mid\mathbf{x}),y\right)$. Specifically, we update SGLD process along the opposite direction of $\nabla_{\mathbf{x}}\log{p_{\theta}}\left((\tilde{\mathbf{x}}\mid\mathbf{x}),y\right)$ to sample the adversarial examples:
$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{c^{2}}{2}\frac{\partial\log{p_{\theta}}\left((\tilde{\mathbf{x}}\mid\mathbf{x}),y\right)}{\partial\tilde{\mathbf{x}}_{t}},$ (12)
[…]
@@
$\frac{\partial{\log}p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\tilde{\mathbf{x}}_{i}|\mathbf{x}_{i}^{+}\right)\right].$ (13)
@@
1Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M$: sampling iterations; $M_{adv}$: adversarial sampling iterations; $M_{\theta}$: sampling iterations for $\theta$ ;
2 Output: $\theta$: network weight;
3 Init: randomly initialize $\theta$;
@@ 7   Obtain ${\mathbf{x}}_{0}$ via random noise;
12  Compute $h_{1}=\frac{\partial\log p_{\theta}({\mathbf{x}})}{\partial\theta}$ via Eq. 9;
@@
14   Obtain $\tilde{\mathbf{x}}_{0}$ from $\mathbf{x}$} with a perturbation;
@@ 4.1 Experimental Setup
We conduct analysis and comparison experiments on several standard classifier datasets, including CIFAR-10 [20], CIFAR-100 [20] and ImageNet [8]. Given that most AT methods utilize WideResNet [43] architectures, we select the commonly used WRN28-10 as the baseline network. All experiments were carried out on 3090 GPU. During the entire training process, we follow the default settings and SGLD hyperparameters of JEMs [39, 41], and the learning rate is set to 0.01. The initialization perturbation for adversarial samples is set to 8/255, and adversarial sampling steps is 5, while the perturbations are constrained using an $\ell_{\infty}$ norm. Note that EB-JDAT is a general optimization framework for JEM that is compatible with various JEM-based models [13, 39, 41, 40]. Therefore, we integrate EB-JDAT with SADAJEM [41] and JEM++ [39], two faster and more stable JEMs, to achieve improved performance.
@@ 4.2 Comparison Experiments And Analysis
We compare our approach with a range of AT methods, including three popular ones: MART [34], AWP [37], and LBGAT [7], as well as six SOTA methods: LAS-AWP [19], UIAT [9], SGLR [23], TDAT [31], AGR-TRADES [30], and DHAT-CFA [45]. For evaluation, we select the PGD-20 [25] and AutoAttack (AA) [6] which consists of APGD-CE [6], APGD-DLR [6], FAB [5] and Square [1]. Both attacks utilizing the $\ell_{\infty}$ norm. All ATs are training on WRN28-10 [43]. The results on CIFAR-10 and CIFAR-100 are shown in Tab. 2, where EB-JDAT not only maintains the classification accuracy of clean data (90.39% on CIFAR-10; 68.35% on CIFAR-100) but also achieves robustness that even surpasses existing SOTA AT methods (+6.38% on CIFAR-10; +1.23% on CIFAR-100). To evaluate the performance of our method on large-scale datasets, we further conducted comparative experiments on the ImageNet subset [8] in Tab. 4. Given that adversarial training on the full ImageNet dataset requires substantial computational resources [38]), we selected a subset consisting of the first 100 categories from ImageNet-1000 and down-sampled the training images to a resolution of $64\times 64$. Thus, our method effectively balances the trade-off between accuracy on clean data and robustness against adversarial attacks. Furthermore, EB-JDAT is compatible with both SADAJEM [41] and JEM++ [39], thereby demonstrating the universality of our approach.
Table 2: Comparative robustness (%) analysis with different AT methods. Results in bold indicate the best.
@@
   Standard Training    96.10    0.06    0    80.73    0    0
   MART [34]    82.99    55.48    50.67    54.69    31.90    27.25
   AWP [37]    82.67    57.21    51.90    57.94    33.75    28.90
   LBGAT [7]    86.22    52.66    50.23    58.64    32.75    28.33
   LAS-AWP [19]    87.74    60.16    55.52    64.89    36.36    30.77
   UIAT [9]    82.94    58.12    52.17    57.65    33.91    29.03
   SGLR [23]    85.72    56.10    53.40    61.02    32.98    28.50
   TDAT [31]    82.25    56.03    54.06    57.32    33.17    26.61
   AGR-TRADES [30]    86.50    52.56    52.56    58.17    28.34    25.84
   DHAT-CFA [45]    84.49    62.38    54.05    61.54    37.15    30.93
   EB-JDAT-JEM++    90.37    64.88    64.61    68.09    37.14    35.21
   EB-JDAT-SADAJEM    90.39    68.76    66.30    68.35    38.38    35.63
Table 3: Camparative robustness (%) analysis with AT methods that utilize generative models for data augmentation. Time is 3090 GPU hours. Results in bold indicate the best.
  Data   Generate data   Epoch   Clean   AA   Time
@@
  SCORE [28]   1M   400   88.10   61.51   $\approx$ 1,438 h
  The method proposed by [12]   100M   2000   87.50   63.38   $\approx$ 719,460 h
  Better DM [35]   1M   400   91.12   63.35   $\approx$ 1,438 h
  EB-JDAT-JEM++   n/a   100   90.37   64.61   31.66 h
  EB-JDAT-SADAJEM   n/a   100   90.39   66.30   66.64 h
@@
  SCORE [28]   1M   400   62.08   31.40   $\approx$ 1,438 h
  Better DM [35]   1M   400   68.06   35.65   $\approx$ 1,438 h
  EB-JDAT-JEM++   n/a   100   68.09   35.21   31.66 h
  EB-JDAT-SADAJEM   n/a   100   68.35   35.63   66.70 h
Table 4: Comparison on ImageNet-Subset in terms of accuracy and robustness. Results in bold indicate the best.
[…]
MART [34] 49.30 21.84 17.40
LAS-AT [19] 50.66 27.34 21.78
WEAT [26] 61.90 30.27 24.52
EB-JDAT-JEM++ 63.02 34.50 32.40
Figure 3: Performance of generated samples on ImageNet subset (64x64) with EB-JDAT-JEM++.
Table 5: Comparison of different hybrid models on CIFAR-10, IF (informative initialization). Results in bold indicate the best.
@@
JEM [13] 92.90 4.28 38.40 8.76
JEM++ [39] 93.73 41.06 37.12 8.29
SADAJEM [41] 96.03 29.63 17.38 8.07
@@
JEAT [46] 85.16 28.43 38.24 8.80
WEAT [26](PCA) 83.36 49.02 30.74 8.97
WEAT [26](IF) 83.36 49.02 177.92 3.50
EB-JDAT-JEM++ 90.37 64.61 39.67 7.66
EB-JDAT-SADAJEM 90.39 66.30 27.42 8.05
Comparisons with AT Using Generated Data.
Another strategy for improving robustness using generative models is to incorporate additional training data generated by generative models in AT [28, 12]. Therefore we compare EB-JDAT with SCORE [28], Better DM [35] and the method proposed in [12]. SOCRE and Better DM utilize the SOTA Diffusion models [15] to generated addition training data, while [12] generate addition data from a ensemble of DDPM [15], GANs [11] and VDVAE [3] generators. All methods are implemented using the WRN28-10 [43]. The results are presented in Tab. 3. Although EB-JDAT does not rely on additional generated data or increased training epochs, it still achieves competitive performance in terms of clean accuracy and robustness. EB-JDAT even outperforms SCORE-1M and [12]. Unlike discriminative learning approaches [28, 12, 35], which struggles to capture the underlying structure of both clean and adversarial data distributions, EB-JDAT, as a hybrid generative model, provides more accurate gradient estimations of clean/adversarial data densities, leading to substantial improvements in both accuracy and robustness. Notably, this is achieved without requiring additional training data and with minimal computational overhead.
Compared with JEMs and Energy-based AT.
In this section, we compare EB-JDAT with other JEMs [13, 39, 41] and energy-based AT [46, 26] in terms of classification accuracy, robustness and generative performance. Tab. 5 presents an evaluation of our method against JEM variants with respect to classification accuracy and generation quality. To assess generation quality, we employ the Fréchet Inception Distance (FID) and Inception Score (IS) as metrics. Compared to JEMs, EB-JDAT maintains comparable accuracy and generative performance, while EB-JDAT-SADAJEM even outperforms JEM (FID-10.98) and JEM++ (FID-9.7) in terms of generation. Moreover, EB-JDAT exhibits significantly improved robustness over JEMs, although it exhibits a slight decrease in classification accuracy and generative ability relative to the strongest JEMs [40], as shown in Tab. 5. This reflects an intrinsic trade-off: adversarial training uses adversarial examples to regularize decision boundaries and reshape the energy landscape, which can shift probability mass away from the clean data manifold and slightly reduce sample fidelity, which are shown in Fig. 2. Overall, compared with existing JEMs [13, 39, 41], EB-JDAT achieves the best balance among classification, generation, and robustness.
Next, we compare EB-JDAT with energy-based AT [46, 26]. First, we compare the accuracy and robustness with them, which is reported in Fig. 1. The results demonstrate that our method significantly enhances robustness with minimal accuracy degradation, outperforming existing energy-based AT. Moreover, our method yields superior generative performance compared to energy‐based methods, which are shown in Tab. 5. EB-JDAT surpasses JEAT and WEAT with an FID reduction of 10.82 and 3.32, respectively. Unlike JEAT [46] directly introduces adversarial samples into the JEMs training process to model the joint distribution ${p}_{\theta}(\tilde{\mathbf{x}},y)$, we model the joint probability $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ to capture the distribution of clean and adversarial samples more accurately, thus performing better in accuracy, robustness and generative ability. On top of that, WEAT [26] reinterprets TRADES [44] as an EBM to model the conditional distribution $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$, which is still a discriminative model. Although [26] notes that PCA-based initialization can give WEAT generative capabilities (FID=30.74), but under the same comparison of informative initialization, WEAT shows poor generative performance (FID=177.92). Consequently, EB-JDAT outperforms previous energy-based AT in classification accuracy, robustness, and generative capability.
@@
To further visually explore generative ability of EB-JDAT, we generate images via SGLD sampling with informative initialization for conventional AT-trained models, energy-based AT-models, JEMs, and EB-JDAT in Fig. 4. As demonstrated in Fig. 4, the conventional AT-trained models (a, b) and energy-based AT-models (e) possess limited generative capacity, resulting in difficulties in generating high-quality images. Conversely, EB-JDAT produce higher-resolution images with richer background details, which is more closely matching the features of the original dataset. In addition, EB-JDAT (g, h) surpasses JEM (f) and perform competitively with JEM++ and SADAJEM in generative capacity, producing more diverse and detailed images with only a slight reduction in sharpness, particularly in the ’car’ category. We also implement our method on the ImageNet subset and generate higher resolution images, as shown in Fig. 3.